The following recipe will make 12 oz of a teriyaki sauce.
Ingredient Amount
Water 1 cup
Soy Sauce 1/4 cup
Brown Sugar 5 tbl
Honey 1 tbl
Ground Ginger 1/2 tsp

How many cups of soy sauce are needed to make 4 oz of teriyaki sauce?
______cups of soy sauce (answer as a reduced fraction)

Found 2 solutions by Theo, math_helper:
Answer by Theo(13342)

12 ounces of teriyaki sauce requires 1/4 cup of soy sauce.
1 cup = 16 tablespoons, therefore:
12 ounces of teriyaki sauce requires 1/4 * 16 = 4 tablespoons of soy sauce.
4 ounces of teriyaki sauce = 4/12 = 1/3 * 12 ounces of teriyaki sauce.
therefore:
4 ounces of teriyaki sauce requires 1/3 * 4 = 4/3 tablespoons of soy sauce.

Answer by math_helper(2461)

Wait, wouldn't it be (1/3)*(1/4 cup) = 1/12 cup? To make 12oz, you need 1/4 cup of soy sauce, so to make 1/3rd as much you'd need 1/3rd as much soy sauce.
---
Using a proportion: 12/4 = (1/4)/x --> x = (1/4)*4/12 = 1/12 cups (the question asks for the answer in cups, not tablespoons)